A boy now one third as old as his father ,twelve years later he will be half as old as his father.then the present age of the boy is
Answer by stanbon(75887)

A boy now one third as old as his father ,twelve years later he will be half as old as his father.then the present age of the boy is
--------
Equations:
b = (1/3)f
b+12 = (1/2)(f+12)
------
Substitute for "b" and solve for "f"::
(1/3)f + 12 = (1/2)(f+12)
-----
Multiply thru by 6 to get rid of the fractions::
2f + 72 = 3(f+12)
2f + 72 = 3f + 36
----
f = 36 (father's age now)
-----
boy = (1/3)f = 12 yrs
------------
Cheers,
Stan H.
----------